OpenStudy (sillybutt100):

Can somebody please teach me how to put sqrt(m^8) in simplified form? The equation will also be in the comments

OpenStudy (luigi0210):

\[\sqrt[8]{m}=m^{\frac{ 8 }{ 2 }}\]

OpenStudy (sillybutt100):

So you take m^8 and divide it by 2?

OpenStudy (luigi0210):

yup..

OpenStudy (sillybutt100):

Thank you!

OpenStudy (sillybutt100):

@Luigi0210 One more question sorry. Do you do the same with ones like this?\[\sqrt{98x ^{5}}\]

OpenStudy (luigi0210):

\[ \sqrt{49*2*x^4*x}\]

OpenStudy (sillybutt100):

So you divide all numbers by two make the variables even and divide those by two?

OpenStudy (luigi0210):

Hm, yea

OpenStudy (sillybutt100):

Thank you again

OpenStudy (luigi0210):

\[7x^2\sqrt{2x}\]

OpenStudy (luigi0210):

Np :D